# Artin Symmetric Functions

### v2

@@ 4.3 Stable Product Identity for Artin L-Functions
Using Brauer’s Theorem 3.7 and the induction property of $L(s,\rho;L/K)$, it suffices to prove this statement for $\dim(\rho)=1.$ In case we will write $\rho=\chi$ for a group character $\chi$ of $\operatorname{\mathrm{Gal}}(L/K)$ and let $f(\chi)\in\mathcal{J}_{K}^{+}$ be the conductor ideal of $\chi.$ We will show that $L^{(n)}(s+n-1,\chi;L/K)$ converges uniformly in compact regions to both the LHS and RHS separately. This will, therefore, show that both LHS and RHS agree in the region $\Re(s)>1.$
@@
$\log h^{(n)}(s)=\sum_{(\operatorname{\mathfrak{q}}^{m},f(\chi))=1}\frac{1}{m}\frac{\chi(\sigma_{\operatorname{\mathfrak{q}}})^{m}}{N(\operatorname{\mathfrak{q}}^{m})^{s}}\frac{1-N(\operatorname{\mathfrak{q}})^{-n}}{1-N(\operatorname{\mathfrak{q}})^{-1}}$
where we sum over all prime power ideals $\operatorname{\mathfrak{q}}^{m}$ in $\mathcal{J}_{K}^{+}$ which are coprime to $f(\chi).$ It suffices to show that for all $\epsilon>0$
$\log h^{(n)}(s)\rightarrow\sum_{(\operatorname{\mathfrak{q}}^{m},f(\chi))=1}\frac{1}{m}\frac{\chi(\sigma_{\operatorname{\mathfrak{q}}})^{m}}{N(\operatorname{\mathfrak{q}}^{m})^{s}}\frac{1}{1-N(\operatorname{\mathfrak{q}})^{-1}}$
@@
$\sum_{(\operatorname{\mathfrak{q}}^{m},f(\chi))=1}\frac{1}{m}\frac{\chi(\sigma_{\operatorname{\mathfrak{q}}})^{m}}{N(\operatorname{\mathfrak{q}}^{m})^{s}}\frac{1}{1-N(\operatorname{\mathfrak{q}})^{-1}}$
@@
$\left|\sum_{(\operatorname{\mathfrak{q}}^{m},f(\chi))=1}\frac{1}{m}\frac{\chi(\sigma_{\operatorname{\mathfrak{q}}})^{m}}{N(\operatorname{\mathfrak{q}}^{m})^{s}}\frac{1}{1-N(\operatorname{\mathfrak{q}})^{-1}}-\sum_{(\operatorname{\mathfrak{q}}^{m},f(\chi))=1}\frac{1}{m}\frac{\chi(\sigma_{\operatorname{\mathfrak{q}}})^{m}}{N(\operatorname{\mathfrak{q}}^{m})^{s}}\frac{1-N(\operatorname{\mathfrak{q}})^{-n}}{1-N(\operatorname{\mathfrak{q}})^{-1}}\right|$
$=\left|\sum_{(\operatorname{\mathfrak{q}}^{m},f(\chi))=1}\frac{1}{m}\frac{\chi(\sigma_{\operatorname{\mathfrak{q}}})^{m}}{N(\operatorname{\mathfrak{q}}^{m})^{s}}\frac{N(\operatorname{\mathfrak{q}})^{-n}}{1-N(\operatorname{\mathfrak{q}})^{-1}}\right|$
$\leq 2\left|\sum_{(\operatorname{\mathfrak{q}}^{m},f(\chi))=1}\frac{1}{m}\frac{N(\operatorname{\mathfrak{q}})^{-n}}{N(\operatorname{\mathfrak{q}}^{m})^{1+\epsilon}}\right|$
$\leq 2^{1-n}\left|\sum_{(\operatorname{\mathfrak{q}}^{m},f(\chi))=1}\frac{1}{m}\frac{1}{N(\operatorname{\mathfrak{q}}^{m})^{1+\epsilon}}\right|$
@@
$\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}^{(n)}}\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{s+n-1}}\operatorname{\mathrm{Tr}}_{\widetilde{\mathbb{H}}_{\lambda^{\bullet}}(W_{\rho})}(\sigma_{K})$
@@
Now using Lemma 4.22, we know for $\rho=\chi$ that if the set of all $\operatorname{\mathfrak{q}}$ with $\lambda^{(\operatorname{\mathfrak{q}})}\neq\emptyset$ is coprime with $f(\chi)$ then $\operatorname{\mathrm{Tr}}_{\widetilde{\mathbb{H}}_{\lambda^{\bullet}}(W_{\rho})}(\sigma_{K})=\prod_{\operatorname{\mathfrak{q}}}\chi(\sigma_{\operatorname{\mathfrak{q}}})^{|\lambda^{(\operatorname{\mathfrak{q}})}|}$ and otherwise $\operatorname{\mathrm{Tr}}_{\widetilde{\mathbb{H}}_{\lambda^{\bullet}}(W_{\rho})}(\sigma_{K})=0.$ Therefore, we can rewrite the sequence in the form
$\sum_{\lambda^{\bullet}\in\mathbb{Y}_{K}}\mathbbm{1}\left(\ell(\lambda^{\bullet})\leq n\right)\varphi_{f(\chi)}(\lambda^{\bullet})\prod_{\operatorname{\mathfrak{q}}}\chi(\sigma_{\operatorname{\mathfrak{q}}})^{|\lambda^{(\operatorname{\mathfrak{q}})}|}\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{n-1}}||\lambda^{\bullet}||^{-s}$
@@
$\left|\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\gamma(\lambda^{\bullet})}{||\lambda^{\bullet}||^{s}}\varphi_{f(\chi)}(\lambda^{\bullet})\prod_{\operatorname{\mathfrak{q}}}\chi(\sigma_{\operatorname{\mathfrak{q}}})^{|\lambda^{(\operatorname{\mathfrak{q}})}|}-\sum_{\lambda^{\bullet}\in\mathbb{Y}_{K}}\mathbbm{1}\left(\ell(\lambda^{\bullet})\leq n\right)\varphi_{f(\chi)}(\lambda^{\bullet})\prod_{\operatorname{\mathfrak{q}}}\chi(\sigma_{\operatorname{\mathfrak{q}}})^{|\lambda^{(\operatorname{\mathfrak{q}})}|}\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{n-1}}||\lambda^{\bullet}||^{-s}\right|$
$=\left|\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\varphi_{f(\chi)}(\lambda^{\bullet})\prod_{\operatorname{\mathfrak{q}}}\chi(\sigma_{\operatorname{\mathfrak{q}}})^{|\lambda^{(\operatorname{\mathfrak{q}})}|}}{||\lambda^{\bullet}||^{s}}\left(\gamma(\lambda^{\bullet})-\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{n-1}}\mathbbm{1}\left(\ell(\lambda^{\bullet})\leq n\right)\right)\right|$
$\leq\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\varphi_{f(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\left|\gamma(\lambda^{\bullet})-\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{n-1}}\mathbbm{1}\left(\ell(\lambda^{\bullet})\leq n\right)\right|$
$=\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\varphi_{f(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\left(\gamma(\lambda^{\bullet})-\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{n-1}}\mathbbm{1}\left(\ell(\lambda^{\bullet})\leq n\right)\right)$
$=\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\varphi_{f(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\gamma(\lambda^{\bullet})-\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\varphi_{f(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{n-1}}\mathbbm{1}\left(\ell(\lambda^{\bullet})\leq n\right)$
@@
$\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\varphi_{f(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{n-1}}\mathbbm{1}\left(\ell(\lambda^{\bullet})\leq n\right)$
@@
$\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\varphi_{f(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\gamma(\lambda^{\bullet}).$
@@
$g_{n,\epsilon}(\lambda^{\bullet}):=\frac{\varphi_{f(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{n-1}}\mathbbm{1}\left(\ell(\lambda^{\bullet})\leq n\right)$
@@
$\widehat{g}_{\epsilon}(\lambda^{\bullet}):=\frac{\varphi_{f(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\gamma(\lambda^{\bullet}).$
@@
$\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\varphi_{f(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\gamma(\lambda^{\bullet})$
@@
$\lim_{n\rightarrow\infty}\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\varphi_{f(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{n-1}}\mathbbm{1}\left(\ell(\lambda^{\bullet})\leq n\right)=\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\varphi_{f(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\gamma(\lambda^{\bullet}).$
@@
$\lim_{n\rightarrow\infty}\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}^{(n)}}\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{s+n-1}}\operatorname{\mathrm{Tr}}_{\widetilde{\mathbb{H}}_{\lambda^{\bullet}}(W_{\rho})}(\sigma_{K})=\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\gamma(\lambda^{\bullet})}{||\lambda^{\bullet}||^{s}}\operatorname{\mathrm{Tr}}_{\widetilde{\mathbb{H}}_{\lambda^{\bullet}}(W_{\rho})}(\sigma_{K})$
@@
$\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\gamma(\lambda^{\bullet})}{||\lambda^{\bullet}||^{s}}\operatorname{\mathrm{Tr}}_{\widetilde{\mathbb{H}}_{\lambda^{\bullet}}(W_{\rho})}(\sigma_{K})$
as desired. ∎

### v3

@@ 4.3 Stable Product Identity for Artin L-Functions
Remark 15.
Note that
$\gamma(\lambda^{\bullet})=\frac{1}{g(\lambda^{\bullet})}P_{\lambda^{\bullet}}[\operatorname{\mathfrak{X}}_{K}]|_{x_{i}^{(\operatorname{\mathfrak{q}})}=N(\operatorname{\mathfrak{q}})^{-(i-1)}}.$
This means that the linear map $\psi:A_{K}\rightarrow\mathbb{Q}^{\text{ab}}$ defined by $\psi(P_{\lambda^{\bullet}}):=g(\lambda^{\bullet})\gamma(\lambda^{\bullet})$ is an algebra homomorphism. This is an arithmetic version of the usual principal specialization on symmetric functions given by $f[\operatorname{\mathfrak{X}}]\rightarrow f\left[\frac{1}{1-t}\right]=f(1,t,t^{2},\ldots).$
@@
We begin by proving this statement for $\dim(\rho)=1.$ To that end we will write $\rho=\chi$ for a group character $\chi$ of $\operatorname{\mathrm{Gal}}(L/K)$ and let $\mathfrak{b}(\chi)\in\mathcal{J}_{K}^{+}$ be the ideal given as the product over all $\mathfrak{q}\in\mathbb{P}_{K}$ such that $\chi|_{I(\mathfrak{q})}$ is nontrivial. We will show that $L^{(n)}(s+n-1,\chi;L/K)$ converges uniformly in compact regions to both the LHS and RHS separately. This will, therefore, show that both LHS and RHS agree in the region $\Re(s)>1.$
@@
$\log h^{(n)}(s)=\sum_{(\operatorname{\mathfrak{q}}^{m},\mathfrak{b}(\chi))=1}\frac{1}{m}\frac{\chi(\sigma_{\operatorname{\mathfrak{q}}})^{m}}{N(\operatorname{\mathfrak{q}}^{m})^{s}}\frac{1-N(\operatorname{\mathfrak{q}})^{-n}}{1-N(\operatorname{\mathfrak{q}})^{-1}}$
where we sum over all prime power ideals $\operatorname{\mathfrak{q}}^{m}$ in $\mathcal{J}_{K}^{+}$ which are coprime to $\mathfrak{b}(\chi).$ It suffices to show that for all $\epsilon>0$
$\log h^{(n)}(s)\rightarrow\sum_{(\operatorname{\mathfrak{q}}^{m},\mathfrak{b}(\chi))=1}\frac{1}{m}\frac{\chi(\sigma_{\operatorname{\mathfrak{q}}})^{m}}{N(\operatorname{\mathfrak{q}}^{m})^{s}}\frac{1}{1-N(\operatorname{\mathfrak{q}})^{-1}}$
@@
$\sum_{(\operatorname{\mathfrak{q}}^{m},\mathfrak{b}(\chi))=1}\frac{1}{m}\frac{\chi(\sigma_{\operatorname{\mathfrak{q}}})^{m}}{N(\operatorname{\mathfrak{q}}^{m})^{s}}\frac{1}{1-N(\operatorname{\mathfrak{q}})^{-1}}$
@@
$\left|\sum_{(\operatorname{\mathfrak{q}}^{m},\mathfrak{b}(\chi))=1}\frac{1}{m}\frac{\chi(\sigma_{\operatorname{\mathfrak{q}}})^{m}}{N(\operatorname{\mathfrak{q}}^{m})^{s}}\frac{1}{1-N(\operatorname{\mathfrak{q}})^{-1}}-\sum_{(\operatorname{\mathfrak{q}}^{m},\mathfrak{b}(\chi))=1}\frac{1}{m}\frac{\chi(\sigma_{\operatorname{\mathfrak{q}}})^{m}}{N(\operatorname{\mathfrak{q}}^{m})^{s}}\frac{1-N(\operatorname{\mathfrak{q}})^{-n}}{1-N(\operatorname{\mathfrak{q}})^{-1}}\right|$
$=\left|\sum_{(\operatorname{\mathfrak{q}}^{m},\mathfrak{b}(\chi))=1}\frac{1}{m}\frac{\chi(\sigma_{\operatorname{\mathfrak{q}}})^{m}}{N(\operatorname{\mathfrak{q}}^{m})^{s}}\frac{N(\operatorname{\mathfrak{q}})^{-n}}{1-N(\operatorname{\mathfrak{q}})^{-1}}\right|$
$\leq 2\left|\sum_{(\operatorname{\mathfrak{q}}^{m},\mathfrak{b}(\chi))=1}\frac{1}{m}\frac{N(\operatorname{\mathfrak{q}})^{-n}}{N(\operatorname{\mathfrak{q}}^{m})^{1+\epsilon}}\right|$
$\leq 2^{1-n}\left|\sum_{(\operatorname{\mathfrak{q}}^{m},\mathfrak{b}(\chi))=1}\frac{1}{m}\frac{1}{N(\operatorname{\mathfrak{q}}^{m})^{1+\epsilon}}\right|$
@@
$\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}^{(n)}}\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{s+n-1}}\operatorname{\mathrm{Tr}}_{\widetilde{\mathbb{H}}_{\lambda^{\bullet}}(W_{\chi})}(\sigma_{K})$
@@
$\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\gamma(\lambda^{\bullet})}{||\lambda^{\bullet}||^{s}}\operatorname{\mathrm{Tr}}_{\widetilde{\mathbb{H}}_{\lambda^{\bullet}}(W_{\chi})}(\sigma_{K}).$
Now using Lemma 4.22, we know for $\rho=\chi$ that if the set of all $\operatorname{\mathfrak{q}}$ with $\lambda^{(\operatorname{\mathfrak{q}})}\neq\emptyset$ is coprime with $\mathfrak{b}(\chi)$ then $\operatorname{\mathrm{Tr}}_{\widetilde{\mathbb{H}}_{\lambda^{\bullet}}(W_{\chi})}(\sigma_{K})=\prod_{\operatorname{\mathfrak{q}}}\chi(\sigma_{\operatorname{\mathfrak{q}}})^{|\lambda^{(\operatorname{\mathfrak{q}})}|}$ and otherwise $\operatorname{\mathrm{Tr}}_{\widetilde{\mathbb{H}}_{\lambda^{\bullet}}(W_{\chi})}(\sigma_{K})=0.$ Therefore, we can rewrite the sequence in the form
$\sum_{\lambda^{\bullet}\in\mathbb{Y}_{K}}\mathbbm{1}\left(\ell(\lambda^{\bullet})\leq n\right)\varphi_{\mathfrak{b}(\chi)}(\lambda^{\bullet})\prod_{\operatorname{\mathfrak{q}}}\chi(\sigma_{\operatorname{\mathfrak{q}}})^{|\lambda^{(\operatorname{\mathfrak{q}})}|}\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{n-1}}||\lambda^{\bullet}||^{-s}$
@@
$\left|\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\gamma(\lambda^{\bullet})}{||\lambda^{\bullet}||^{s}}\varphi_{\mathfrak{b}(\chi)}(\lambda^{\bullet})\prod_{\operatorname{\mathfrak{q}}}\chi(\sigma_{\operatorname{\mathfrak{q}}})^{|\lambda^{(\operatorname{\mathfrak{q}})}|}-\sum_{\lambda^{\bullet}\in\mathbb{Y}_{K}}\mathbbm{1}\left(\ell(\lambda^{\bullet})\leq n\right)\varphi_{\mathfrak{b}(\chi)}(\lambda^{\bullet})\prod_{\operatorname{\mathfrak{q}}}\chi(\sigma_{\operatorname{\mathfrak{q}}})^{|\lambda^{(\operatorname{\mathfrak{q}})}|}\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{n-1}}||\lambda^{\bullet}||^{-s}\right|$
$=\left|\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\varphi_{\mathfrak{b}(\chi)}(\lambda^{\bullet})\prod_{\operatorname{\mathfrak{q}}}\chi(\sigma_{\operatorname{\mathfrak{q}}})^{|\lambda^{(\operatorname{\mathfrak{q}})}|}}{||\lambda^{\bullet}||^{s}}\left(\gamma(\lambda^{\bullet})-\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{n-1}}\mathbbm{1}\left(\ell(\lambda^{\bullet})\leq n\right)\right)\right|$
$\leq\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\varphi_{\mathfrak{b}(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\left|\gamma(\lambda^{\bullet})-\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{n-1}}\mathbbm{1}\left(\ell(\lambda^{\bullet})\leq n\right)\right|$
$=\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\varphi_{\mathfrak{b}(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\left(\gamma(\lambda^{\bullet})-\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{n-1}}\mathbbm{1}\left(\ell(\lambda^{\bullet})\leq n\right)\right)$
$=\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\varphi_{\mathfrak{b}(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\gamma(\lambda^{\bullet})-\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\varphi_{\mathfrak{b}(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{n-1}}\mathbbm{1}\left(\ell(\lambda^{\bullet})\leq n\right)$
@@
$\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\varphi_{\mathfrak{b}(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{n-1}}\mathbbm{1}\left(\ell(\lambda^{\bullet})\leq n\right)$
@@
$\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\varphi_{\mathfrak{b}(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\gamma(\lambda^{\bullet}).$
@@
$g_{n,\epsilon}(\lambda^{\bullet}):=\frac{\varphi_{\mathfrak{b}(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{n-1}}\mathbbm{1}\left(\ell(\lambda^{\bullet})\leq n\right)$
@@
$\widehat{g}_{\epsilon}(\lambda^{\bullet}):=\frac{\varphi_{\mathfrak{b}(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\gamma(\lambda^{\bullet}).$
@@
$\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\varphi_{\mathfrak{b}(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\gamma(\lambda^{\bullet})$
@@
$\lim_{n\rightarrow\infty}\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\varphi_{\mathfrak{b}(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{n-1}}\mathbbm{1}\left(\ell(\lambda^{\bullet})\leq n\right)=\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\varphi_{\mathfrak{b}(\chi)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{1+\epsilon}}\gamma(\lambda^{\bullet}).$
@@
$\lim_{n\rightarrow\infty}\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}^{(n)}}\frac{\kappa^{(n)}(\lambda^{\bullet})}{||\lambda^{\bullet}||^{s+n-1}}\operatorname{\mathrm{Tr}}_{\widetilde{\mathbb{H}}_{\lambda^{\bullet}}(W_{\chi})}(\sigma_{K})=\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\gamma(\lambda^{\bullet})}{||\lambda^{\bullet}||^{s}}\operatorname{\mathrm{Tr}}_{\widetilde{\mathbb{H}}_{\lambda^{\bullet}}(W_{\chi})}(\sigma_{K})$
@@
$\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\gamma(\lambda^{\bullet})}{||\lambda^{\bullet}||^{s}}\operatorname{\mathrm{Tr}}_{\widetilde{\mathbb{H}}_{\lambda^{\bullet}}(W_{\chi})}(\sigma_{K})$
as desired.
Now suppose $\rho$ is arbitrary. Using Brauer’s theorem 3.7 we may write
$L(s,\rho;L/K)=\prod_{i=1}^{r}L(s,\operatorname{\mathrm{Ind}}(\chi_{i});L/K)^{a_{i}}$
where $\chi_{i}:\operatorname{\mathrm{Gal}}(L/K_{i})\rightarrow\mathbb{C}^{*}$ are $1$-dimensional characters, $L/K_{i}/K$ are some intermediate extensions, $\operatorname{\mathrm{Ind}}(\chi_{i})$ is shorthand for $\operatorname{\mathrm{Ind}}_{\operatorname{\mathrm{Gal}}(L/K_{i})}^{\operatorname{\mathrm{Gal}}(L/K)}(\chi)$, and $a_{i}\in\mathbb{Z}.$ For all $n\geq 0$ we have that
$\prod_{j\geq 0}^{n}L(s+j,\rho;L/K)=\prod_{i=1}^{r}\left(\prod_{j\geq 0}^{n}L(s+j,\operatorname{\mathrm{Ind}}(\chi_{i});L/K)\right)^{a_{i}}.$
But now we have that
$\prod_{j\geq 0}^{n}L(s+j,\operatorname{\mathrm{Ind}}(\chi_{i});L/K)=\prod_{j\geq 0}^{n}L(s+j,\chi_{i};L/K_{i})$
so we may apply our prior computation to obtain that for all $1\leq i\leq r,$
$\lim_{n\rightarrow\infty}\prod_{j\geq 0}^{n}L(s+j,\operatorname{\mathrm{Ind}}(\chi_{i});L/K)=\prod_{j\geq 0}L(s+j,\operatorname{\mathrm{Ind}}(\chi_{i});L/K)$
converges uniformly in compact subsets of $\Re(s)>1$ to an analytic function. But then we know that
$\lim_{n\rightarrow\infty}\prod_{j\geq 0}^{n}L(s+j,\rho;L/K)=\prod_{j\geq 0}L(s+j,\rho;L/K)$
must also converge uniformly in compact subsets of $\Re(s)>1$ to an analytic function. Similarly, we know that each of the series for $1\leq i\leq r$
$\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K_{i}}}\frac{\gamma(\lambda^{\bullet})}{||\lambda^{\bullet}||^{s}}\operatorname{\mathrm{Tr}}_{\widetilde{\mathbb{H}}_{\lambda^{\bullet}}(W_{\chi_{i}})}(\sigma_{K_{i}})$
defines an analytic function on $\Re(s)>1$ agreeing with $\prod_{j\geq 0}L(s+j,\chi_{i};L/K_{i}).$ By using Theorem 3.23, Proposition 3.4, and the algebra homomorphism $\psi$ from Remark 15 we see that for all $1\leq i\leq r$
$\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K_{i}}}\frac{\gamma(\lambda^{\bullet})}{||\lambda^{\bullet}||^{s}}\operatorname{\mathrm{Tr}}_{\widetilde{\mathbb{H}}_{\lambda^{\bullet}}(W_{\chi_{i}})}(\sigma_{K_{i}})=\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\gamma(\lambda^{\bullet})}{||\lambda^{\bullet}||^{s}}\operatorname{\mathrm{Tr}}_{\widetilde{\mathbb{H}}_{\lambda^{\bullet}}(W_{\operatorname{\mathrm{Ind}}(\chi_{i})})}(\sigma_{K}).$
Now we have that
$\prod_{j\geq 0}L(s+j,\rho;L/K)=\prod_{i=1}^{r}\left(\sum_{\lambda^{\bullet}\in\operatorname{\mathbb{Y}}_{K}}\frac{\gamma(\lambda^{\bullet})}{||\lambda^{\bullet}||^{s}}\operatorname{\mathrm{Tr}}_{\widetilde{\mathbb{H}}_{\lambda^{\bullet}}(W_{\operatorname{\mathrm{Ind}}(\chi_{i})})}(\sigma_{K})\right)^{a_{i}}$
and so the RHS must yield a Dirichlet series which converges uniformly for $\Re(s)>1$ to an analytic function. But again applying Theorem 3.23 and the algebra homomorphism $\psi$ from Remark 15 we see that the coefficients must be given as